MICROWAVE SPECTROSCOPY OF 1,1-DIFLUOROCYANOMETHYL RADICAL, ĊF_{2}CN

\underline{LU KANG}, Department of Natural Sciences, Union College, Barbourville, KY 40906; STEWART E. NOVICK, Department of Chemistry, Wesleyan University, Middletown, CT 06459.

The pure rotational spectrum of 1,1-difluorocyanomethyl radical, ĊF_{2}CN, was measured between 7.0 and 26.5 GHz using a Fabry-Pérot cavity Fourier transform microwave (FP-FTMW) spectrometer. The a-type (K_{a} = 0, 1, 2) paramagnetic transitions were observed and assigned. The spectroscopic constants of ĊF_{2}CN were fit by Pickett^{,}s SPFIT suite of programs. The experimentally determined rotational constants are: A_{0} = 11010.58(1) MHz, B_{0} = 4081.6917(1) MHz, and C_{0} = 2989.9589(1) MHz. These constants are in a good agreement with the predicted values from density functional theory (DFT) calculation, UB3LYP/aug-cc-pVQZ. The inertial defect, Δ = -0.69 amuÅ^{2}, indicates that ĊF_{2}CN has a non-planar geometry. This is not the case in its isoelectronic analog, 1,1-difluoropropargyl, ĊF_{2}CCH ^{a}.

aL. Kang, S. E. Novick; J. Chem. Phys. 125, 054309, (2006)